OpenStudy (anonymous):

Find the linerization of a suitable function and then use it to aproximate the number: (i)1.002^3 (ii)63.8^(1/2) (iii) 31.08^(1/5) (iv) sin 0.1 teach showed me an example but it wasn't like this

OpenStudy (anonymous):

Linearization is L(x)=y(x)+y'(x)dx

OpenStudy (anonymous):

i)1.002^3 Something cube Let say x^3 Dy/dx (x^3) =3x^2

OpenStudy (anonymous):

L(x)=y(x)+y'(x)dx y(x)=x^3 y'(x)=3x^2

OpenStudy (anonymous):

Let's plug them in x^3+3x^2 dx

OpenStudy (anonymous):

any question so far?

OpenStudy (anonymous):

yea that use a suitable function part, so all thats asking is use that L(x) formula?

OpenStudy (anonymous):

i ask because in the example she game us a f(x) formula with it

OpenStudy (anonymous):

and a point of x

OpenStudy (anonymous):

So " use a suitable function part" Some example 2^3 -- use this function x^3 Ln(3) -- use this func ln(x)

myininaya (myininaya):

y=(32-x)^(1/5) y'=-1/5(32-x)^(-4/5) y'=-1/5(32-0)^(-4/5)=-1/5(32)^(-4/5)=-1/5(1/16)=-1/80 y=mx+b y=-1/80*x+b (we know a point on the lint (0,2) y=-1/80*x+2 L(x)=mx+b=-1/80*x+2 so we want to find L at x=.92=92/100=23/25 L(23/25)=-1/80*23/25+2=3977/2000=1.9885 if we have done y=(32-.92)^(1/5)=(31.08)^(1/5)=1.988365 approximately all you are doing is finding the tangent line

OpenStudy (anonymous):

wow, lol gimmie a min to try and digest this, u guys are good,

OpenStudy (anonymous):

wow, didn't even figure all that formula was just for the tangent line, with all the f' of tyhis and that

OpenStudy (anonymous):

thnx alot for the help, its much appreciated guys